OpenStudy (anonymous):

solve for t 5t+12=4-3t prove solution is correct

OpenStudy (jamesj):

First of all, add 3t to both sides: 3t + 5t + 12 = 4 - 3t + 3t 8t + 12 = 4 + 0 Now subtract 12 from both sides and you get 8t = ... what? and hence t = .... what?